### v1

@@ 1.2 Our Contributions
In this paper, we introduce several new results that relax some of the previous assumptions considered in the literature [16] , and, additionally, we introduce new results and computations in the context of general $N$-player $N$-player oligopoly markets, such as those studied in [5]. Furthermore, we exploit the structure of the oligopoly market to derive sharper stability results and characterizations that quantify the influence of deception in this context. We also provide a broader estimate for the stability-preserving set in the nominal average dynamics and offer results that determine when the new equilibrium point under deception is truly a Nash equilibrium, rather than just an equilibrium, for the deceptive game. This framework further allows us to consider a more general class of NES dynamics, where players are permitted to use different gains and amplitudes in their update laws. The effectiveness of these results are also demonstrated through numerical simulations.
@@ 2.1 NES for the oligopoly
Consider an $N$-player noncooperative game, where player $i$ implements action $x_{i}\in\mathbb{R}$ and aims to unilaterally minimize their cost function $J_{i}:\mathbb{R}^{N}\to\mathbb{R}$. We use $[N]:=\{1,2,...,N\}$ to denote the set of players, and we let $x=[x_{1},...,x_{N}]^{\top}$ denote the vector of players’ actions. Similarly we use $x_{-i}\in\mathbb{R}^{N-1}$ to denote the vector of all players’ actions except for the action of player $i$. Given real-valued cost functions $J_{i}(x_{i},x_{-i}):\mathbb{R}^{N}\to\mathbb{R}$, for all $i$, a policy $x^{*}\in\mathbb{R}^{N}$ is called a Nash equilibrium if it satisfies
@@
$[Q_{i}]_{jk}$ =$=\begin{dcases}\frac{2R_{\|}}{R_{i}\overline{R}_{i}},\quad&\text{if }j=k=i\\ -\frac{R_{\|}}{R_{j}R_{k}},\quad&\text{if }j\neq k\text{ and either }j=i\text{or }k=i\\ 0,\quad&\text{else}\end{dcases}$ (7a)
$[b_{i}]_{k}$ =$=\begin{dcases}-\frac{m_{i}R_{\|}}{R_{i}\overline{R}_{i}}-\frac{S_{d}R_{\|}}{R_{i}},\quad&\text{if }k=i\\ \frac{m_{i}R_{\|}}{R_{i}R_{k}},\quad&\text{else}\end{dcases}$ (7b)
Here, $[Q_{i}]_{jk}$ and $[b_{i}]_{k}$ denote the $(j,k)$ entry of $Q_{i}$ and $k$-th entry of $b_{i}$ respectively. It follows immediately that the pseudogradient of the oligopoly can be represented in the form $\mathcal{G}(x)=\mathcal{Q}x+\mathcal{B}$, where $\mathcal{Q}\in\mathcal{R}^{N\times N}$ and $\mathcal{B}\in\mathbb{R}^{N}$ satisfy $[\mathcal{Q}]_{i:}=[Q_{i}]_{i:}$ and $[\mathcal{B}]_{i}=[b_{i}]_{i}$ for all $i$. Here, $[E]_{i:}$ denotes row $i$ of matrix $E$.
@@ 2.2 Deception in the Oligopoly
$x_{i}$ =$=\begin{dcases}u_{i}+a_{i}\sin(\omega_{i}t)+\delta_{i}\sum_{j\in\mathcal{D}_{i}}a_{j}\sin\left(\omega_{j}t\right)&\text{if }i\in\mathcal{D}\\ u_{i}+a_{i}\sin(\omega_{i}t)&\text{else}\end{dcases}$ (10a)
@@ 3.1 Stability
[¯$[\overline{\mathcal{B}}(\delta)]_{i}$ $=[\mathcal{B}]_{i}+\sum_{k\in\mathcal{K}_{i}}\delta_{k}\frac{m_{i}R_{\|}}{R_{i}R_{k}}.$ (11b)
@@
By (7a) we already know $[Q_{i}]_{kj}=0$ for $k\in\mathcal{K}_{i}$ and $j\neq i$, thus for fixed $i$ we have
@@
    The matrix $\Lambda(\delta^{*})\in\mathbb{R}^{n\times n}$ with $[\Lambda(\delta^{*})]_{jk}=\nabla_{j}\xi_{k}(\delta^{*})$ is Hurwitz, where $\xi_{k}:\mathbb{R}^{n}\to\mathbb{R}$ is given by $\xi_{k}(\delta):=\varepsilon_{z_{k}}J_{z_{k}}(-\overline{\mathcal{Q}}(\delta)^{-1}\overline{\mathcal{B}}(\delta))$.
@@
[…]
@@
$\frac{\partial\tilde{u}}{\partial\tau}=\frac{1}{\omega}\left(-K\overline{\mathcal{Q}}(\tilde{\delta})\tilde{u}-K\overline{\mathcal{B}}(\tilde{\delta})\right).$ (17)
@@
where $P_{i}:\mathbb{R}^{N}\to\mathbb{R}$ is a quadratic function satisfying $\mathcal{P}_{i}(a)=0$, and can thus be treated as an $\mathcal{O}(a)$ perturbation on compact sets, where $a=[a_{1},...,a_{N}]^{\top}$. We recall that $\mathcal{D}=\{z_{1},...,z_{n}\}$. By setting $\tau^{*}=\varepsilon\tau$ and ${J}^{*}(\tilde{u})=[J_{z_{1}}(\tilde{u}),...,J_{z_{n}}(\tilde{u})]^{\top}$, the entire system can be represented as
@@
If we disregard the perturbation in (20), the resulting system is a singularly perturbed system with quasi steady state $h(\tilde{\delta})=-\overline{\mathcal{Q}}(\tilde{\delta})^{-1}\overline{\mathcal{B}}(\tilde{\delta})$ and reduced dynamics given by
@@
Since $J^{\text{ref}}\in\Omega$, it follows that (21) has an exponentially stable equilibrium point $\delta^{*}\in\Delta$. Moreover, by denoting $y=\tilde{u}-h(\tilde{\delta})$, we obtain the boundary layer system
@@ 3.2 Intuition Behind Deception
However, note that $[\overline{\mathcal{Q}}(\delta)]_{ij}=[Q_{i}]_{ij}$ for $i\neq j$, and we also have
@@
[…]
@@ 4 Numerical Results
¯$\overline{\mathcal{B}}(\delta)$ $=\begin{bmatrix}-48.82\\ -90.34\\ -51.65\end{bmatrix}+\delta\begin{bmatrix}0\\ 0\\ 10.14\end{bmatrix}$ (33b)

### v2

@@ 1.2 Our Contributions
In this paper, we introduce several new results that relax some of the previous assumptions considered in the literature [16] , and, additionally, we introduce new results and computations in the context of general $N$-player NN-player oligopoly markets, such as those studied in [5]. Furthermore, we exploit the structure of the oligopoly market to derive sharper stability results and characterizations that quantify the influence of deception in this context. We also provide a broader estimate for the stability-preserving set in the nominal average dynamics and offer results that determine when the new equilibrium point under deception is truly a Nash equilibrium, rather than just an equilibrium, for the deceptive game. This framework further allows us to consider a more general class of NES dynamics, where players are permitted to use different gains and amplitudes in their update laws. The effectiveness of these results are also demonstrated through numerical simulations.
@@ 2.1 NES for the oligopoly
Consider an $N$-player noncooperative game, where player $i$ implements action $x_{i}\in\mathbb{R}$ and aims to unilaterally minimize their cost function $J_{i}:\mathbb{R}^{N}\to\mathbb{R}$. We use $N]:=\{1,2,...,N\}$ to denote the set of players, and we let $x=[x_{1},...,x_{N}]^{\top}$ denote the vector of players’ actions. Similarly we use $x_{-i}\in\mathbb{R}^{N-1}$ to denote the vector of all players’ actions except for the action of player $i$. Given real-valued cost functions $J_{i}(x_{i},x_{-i}):\mathbb{R}^{N}\to\mathbb{R}$, for all $i$, a policy $x^{*}\in\mathbb{R}^{N}$ is called a Nash equilibrium if it satisfies
@@
$[Q_{i}]_{jk}$ $=\begin{dcases}\frac{2R_{\|}}{R_{i}\overline{R}_{i}},\quad&\text{if }j=k=i\\ -\frac{R_{\|}}{R_{j}R_{k}},\quad&\text{if }j\neq k\text{ and either }j=i\text{ or }k=i\\ 0,\quad&\text{else}\end{dcases}$ (7a)
$[b_{i}]_{k}$ $=\begin{dcases}-\frac{m_{i}R_{\|}}{R_{i}\overline{R}_{i}}-\frac{S_{d}R_{\|}}{R_{i}},\quad&\text{if }k=i\\ \frac{m_{i}R_{\|}}{R_{i}R_{k}},\quad&\text{else}\end{dcases}$ (7b)
Here, $Q_{i}]_{jk}$ and $b_{i}]_{k}$ denote the $(j,k)$ entry of $Q_{i}$ and $k$-th entry of $b_{i}$ respectively. It follows immediately that the pseudogradient of the oligopoly can be represented in the form $\mathcal{G}(x)=\mathcal{Q}x+\mathcal{B}$, where $\mathcal{Q}\in\mathcal{R}^{N\times N}$ and $\mathcal{B}\in\mathbb{R}^{N}$ satisfy $\mathcal{Q}]_{i:}=[Q_{i}]_{i:}$ and $\mathcal{B}]_{i}=[b_{i}]_{i}$ for all $i$. Here, $E]_{i:}$ denotes row $i$ of matrix $E$.
@@ 2.2 Deception in the Oligopoly
$x_{i}$ $=\begin{dcases}u_{i}+a_{i}\sin(\omega_{i}t)+\delta_{i}\sum_{j\in\mathcal{D}_{i}}a_{j}\sin\left(\omega_{j}t\right)&\text{if }i\in\mathcal{D}\\ u_{i}+a_{i}\sin(\omega_{i}t)&\text{else}\end{dcases}$ (10a)
@@ 3.1 Stability
$[\overline{\mathcal{B}}(\delta)]_{i}$ $=[\mathcal{B}]_{i}+\sum_{k\in\mathcal{K}_{i}}\delta_{k}\frac{m_{i}R_{\|}}{R_{i}R_{k}}.$ (11b)
@@
By (7a) we already know $Q_{i}]_{kj}=0$ for $k\in\mathcal{K}_{i}$ and $j\neq i$, thus for fixed $i$ we have
@@
    The matrix $\Lambda(\delta^{*})\in\mathbb{R}^{n\times n}$ with $\Lambda(\delta^{*})]_{jk}=\nabla_{j}\xi_{k}(\delta^{*})$ is Hurwitz, where $\xi_{k}:\mathbb{R}^{n}\to\mathbb{R}$ is given by $\xi_{k}(\delta):=\varepsilon_{z_{k}}J_{z_{k}}(-\overline{\mathcal{Q}}(\delta)^{-1}\overline{\mathcal{B}}(\delta))$.
@@
$=-\frac{2k_{i}}{a_{i}\omega T}\int_{0}^{T}\sin(\overline{\omega}_{i}\tau)\tilde{\mu}(\tau)^{\top}(Q_{i}\tilde{u}+b_{i})d\tau$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.2 Intuition Behind Deception
However, note that $\overline{\mathcal{Q}}(\delta)]_{ij}=[Q_{i}]_{ij}$ for $i\neq j$, and we also have
@@
Consider the $N$-player oligopoly with costs $J_{i}$ of the form (2). Let $\delta\in\{\delta\in\mathbb{R}^{n}:0<|\delta|<2\emph{ and }\overline{\mathcal{Q}}(\delta)\emph{ is invertible}\}$ and let $\{\tilde{J}_{i}\}_{i\in[N]}$ be a corresponding deceptive game. Then, the point $u^{*}=-\overline{\mathcal{Q}}(\delta)^{-1}\overline{\mathcal{B}}(\delta)$ is a Nash equilibrium of the deceptive game $\{\tilde{J}_{i}\}_{i\in[N]}$.
@@ 4 Numerical Results
$\overline{\mathcal{B}}(\delta)$ $=\begin{bmatrix}-48.82\\ -90.34\\ -51.65\end{bmatrix}+\delta\begin{bmatrix}0\\ 0\\ 10.14\end{bmatrix}$ (33b)